Hi to all, please help me solve these following equations with solution:
How many possible zeros of the given equation x^4- 2x^3 - 8x^2 +12x + 16 = 0?
a. ) 2 b.) 3 c. ) 4 d.) 5
What are the possible zeros of the equation x4 - 2x^3 - 3x^2 + 4x + 4 = 0
a. ) ( x - 2)^2 ( x +1 )^2 b.) ( x +3 )^2 ( x +1 )^2
c. ) ( x +1 )^( x +2 )^2 d.) ( x - 2 )^2 ( x - 2 )^2

Answer by ewatrrr(24785)


Hi
How many possible zeros of the given equation x^4- 2x^3 - 8x^2 +12x + 16 = 0?
4th Degree Equation: d) 4 possible solutions.
Solving or Using a Solver... the result would be 2 real and 2 imaginary roots
What are the possible zeros of the equation x^4 - 2x^3 - 3x^2 + 4x + 4 = 0
a.) ( x - 2)^2 ( x +1 )^2
Multiply:
(x^2 - 4x + 4)
(x^2 +2x + 1) = x^4 - 2x^3 - 3x^2 + 4x + 4
Or one can verify with an online 4th Degree Equation Solver
Wish You the Best in your Studies.